• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name G.M. Urciuoli

    Log entry time 07:59:19 on July24,2003

    Entry number 106441

    keyword=7/24/2003 owl sfit report




    7/24/2003

    Owl shift

    Shift crew:

    SL G.M. Urciuoli
    TO: S. Woo
    3rd:: J. Niedziela


    00:00 owl shift starts; beam off.

    00:30 1.1 uA



    00:32 2722 1.127 empty 5 12 1.2 M

    beam 0.15 uA

    1:00 2723 junk

    1:04 2724 1.116 multifoils 2 12 1.2 M

    1:18 2725 1.116 single 2 12 1. M


    beam 1 uA


    1:32 2726 1.094 multifoils 18 8 1.2 M

    1:48 2727 1.094 single 14 7 junk

    1:49 2728 1.094 single 14 7 1.2 M


    2:00 Septum starts malfunctioning. It is not able to get 99.26 A. The constant Coil Voltage Alarm went off, and the Septum Voltage started to drop to zero as soon as 99.26 A are reached. We tried resetting septum twice (once with beam on, once with beam off) without success. Mark Steven has been callled to fix the problem.

    raster check out: run 2729-2731
    Raster Size 3x3 mm

    3:45 HallA in "controlled access" status

    6:40 the Septum problem is not complety fixed. However, because its current is now
    at the desired value, we continue taking data.

    6:47 1uA beam in Hall.

    Time Run Number mom. target Ps1 Ps2 N. Events

    6:48 2732 1.072 multifoils 23 9 raster on

    raster was accidentaly left on in run 2732.

    7:06 2733 1.072 multifoils 23 9 1.2 M
    7:24 2734 1.072 single 14 5 1.2 M

    when setting septum at 97.23, the magnet tripped again. After it ramped down almost to zero, we were able to set the desired value.

    7:54 2735 junk